OpenStudy (anonymous):

Let setM = {x, 2x, 4x} for any number x. If the average (arithmetic mean) of the numbers in set M is 14, what is the value of x ? a.6 b.7 c.10 d.12 e.14

OpenStudy (tkhunny):

How do you calculate and average?

OpenStudy (anonymous):

average a,b,c,d = (a+b+c+d)/4 in this case we have x, 2x, 4x whose average is 14 this implies that (x+2x+4x)/3 =14 or (x+2x+4x) = (14)(3) or 7x = (14)(3) or x= (14)(3)/7 or x= (2)(3) or x=6